A recipe calls for 2/5 cup of milk. How much milk should be used to make 1/4 of the recipe?
Answer by nycsharkman(136)

A recipe calls for 2/5 cup of milk. How much milk should be used to make 1/4 of the recipe?
1/(1/4) = 2/5 cup of milk/x....This means: 1 divided by 1/4 = 2/5 divided by x.
Let x = the amount of cup of milk
Cross-multiply and solve for x.
1/4 times 2/5 = 1/10
1 times x = x
Final answer: x = 1/10 cup of milk should be used to make 1/4 of the recipe.